OpenStudy (anonymous):

Write an equation of a line perpendicular to y =1/4 x + 4 in slope-intercept form that passes through the point (-2, 6).

OpenStudy (mrnood):

from the equation y=1/4 x +4 what is the slope of that line? (hint: In the form y=mx+b : m is the slope)

OpenStudy (anonymous):

the slope is 1/4 so the new slope would be -4

OpenStudy (anonymous):

im having trouble on how to find the b of the new equation

OpenStudy (anonymous):

@MrNood do you know how to do that ?

OpenStudy (mrnood):

ok so you have the new slope so the equation of th enew line would be y = -4x + b but you know that the point (-2,6) lies on the line so when x=-2 y= 6 so put those 2 numbers in the equation = and you will find b

OpenStudy (anonymous):

okay so 6 = -4 * -2 + b 6 = 8 + b b = -2 Is that right @MrNood

OpenStudy (mrnood):

check it by putting it back into the equation .... (it looks right to me...)

OpenStudy (anonymous):

so the final answer would be y = -4 x - 2 right?

OpenStudy (mrnood):